OpenStudy (mellamokatie):

6. Factor the following polynomial: 12y^5-8y^3+24y^2 A. 8y(2y^4-y^2+3y) B. 4y^2(3y^3-2y+6) C. 4y(3y^5-2y^3+6y^2) D. 4y(3y^4-2y^2_6y) @welshfella

OpenStudy (dopeboydee):

ummmmm.

OpenStudy (mellamokatie):

Hahahaha

OpenStudy (sooobored):

are you familiar with prime factors? what prime factors do each of the terms share?

OpenStudy (mellamokatie):

Are you talking about like terms?

OpenStudy (sooobored):

no i mean ify ou have the term 12xy^2 the prime factors of this would be 2,2,3,x,y,y

OpenStudy (sooobored):

breaking it up into its smallest components, which in terms of numbers, would be prime numbers

OpenStudy (mellamokatie):

I dont know that

OpenStudy (sooobored):

ok, then use division rules any even number can be divided by 2 right?

OpenStudy (dopeboydee):

Its 1 answer down from A. :)

OpenStudy (mellamokatie):

Yes

OpenStudy (sooobored):

so if you look at all your terms, 12y^5, -8y^3, and 24y^2 are all these terms even?

OpenStudy (mellamokatie):

Uh no

OpenStudy (welshfella):

you need to find the GCF of the 3 terms do you know what the GCF is?

OpenStudy (mellamokatie):

No

OpenStudy (welshfella):

the greatest common factor First consider the numbers GCF of 12,8 and 24 is 4 because its the largest number which divides into all 3 numbers

OpenStudy (welshfella):

Next you nee to find the GCF of y^5 , y^3 and y^2

OpenStudy (welshfella):

one of those will divide exactly into all 3 Which one?

OpenStudy (mellamokatie):

D: I dont know

OpenStudy (welshfella):

do you think it might be y^2? y^2 /y^2 = 1 will y^2 divide into y^3 and y^5?

OpenStudy (welshfella):

recall when you divide exponential terms you subtract the exponents

OpenStudy (mellamokatie):

I would think its y^5

OpenStudy (welshfella):

No the GCF is not y^5 because it is not a factor of y^2 or y^3 - but y^2 is a factor of both y^5 /y^2 = y^3 y^3 / y^2 = y so y^2 is the GCF

OpenStudy (mellamokatie):

Oh I see

OpenStudy (welshfella):

combining this with the 4 we have 4y^2 which is the GCF of the 3 terms in the original question

OpenStudy (welshfella):

we can also see that y^2 is the GCF if we expand the terms y^5 = y * y* y * y * y y^3 = y * y * y y^2 = y * y We see that y^2 will divide into all 3.

OpenStudy (mellamokatie):

Ohh okay I like how you arranged it